dx=(10y+6z)dy+6y, dx=(3y2+4yz)dy+(2yz+y2)dz, dx=y4-1dy+zdz


1
Expert's answer
2020-10-20T17:37:46-0400

given,

"dx=(10y+6z)dy+6y"

Integrate both side we get,

"x=5y^2+6yz+6y+c",


dx="(3y^2+4yz)dy+(2yz+y^2)dz"


This is of the form

Mdy+Ndz

so the solution is,

Integrate both the sides,

"x=y^3+2zy^2+c" ,


dx="(y^4-1)dy+zdz"


Integrate both the sides,

"x=\frac{y^5}{5}-y+\frac{z^2}{2}+c"